### v1

@@ Appendix D Validation of the NLTE parameters and abundances
We compared our NLTE atmospheric parameters and chemical abundances with previous studies that investigated the impact of NLTE analysis, focusing in particular on the behavior of NLTE-LTE differences as a function of metallicity. ruchti13 introduced an optimized methodology, termed NLTE-Opt, specifically designed to mitigate the systematic biases typical of standard 1D LTE analyses for FGK-type stars. In this approach, is determined from Balmer lines profiles, while and [Fe/H] are derived from the NLTE ionization balance of Fe. Fig. 29 shows the systematic offsets in stellar parameters as a function of metallicity based on standard LTE analysis when compared with NLTE-Opt estimates. These trends are generally attributed to the inability of the LTE Fe I excitation balance to provide an accurate scale. As shown by the black circles representing the results of ruchti13, these differences decrease moving toward higher metallicities (see also kovalev19). Our Cepheids (in blue) follow the same general behavior, displaying a good agreement with ruchti13, particularly in (panel a), (panel b) and (panel d). However, the panel c) of the same figure shows the difference in microturbulence differences, revealing an offset of 0.20–0.30 . Since Cepheids typically exhibit higher micro- and macroturbulent velocities, both the observed trends and the intrinsic scatter should be interpreted with caution. We also note that is the only parameter for which we find a statistically significant offset between the NLTE and LTE determinations (-0.150.11 , see Fig. 2). Finally, our fitting method simultaneously optimizes all atmospheric parameters, which means that the contribution of significantly affects the determination of the others.
@@
We independently verified our NLTE abundances by applying the line-by-line NLTE abundance corrections provided by the MPIA database101010https://nlte.mpia.de/ (kovalev18). Specifically, we focused on Mn and O. The atomic model for Mn is the same as that used in our study from bergemann19, while for O, we used the atomic model from amarsi2018, while the MPIA database relies on bergemann21. We performed this comparison for five sample spectra whose atmospheric parameters are representative of the whole population. We downloaded the NLTE abundance correction from the MPIA database for Mn (5394.68 , 5420.33 and 6021.77 ) and O (7771.94 , 7774.17 and 7775.39 ). As shown in Fig. 30, the comparison indicates good agreement for all Cepheids, especially for Mn. For oxygen the agreement is also good (differences within 0.1 dex) for three Cepheids, while the remaining two show discrepancies of approximately 0.2 dex. This is well within our observational uncertainties. Finally, the decreasing trend observed in the plane vs has also been obtained in vasilyev19 for the 777.4 nm line with both 1D and 2D models, and previously observed in Cepheids in luck13.

### v2

@@ Appendix D Validation of the NLTE parameters and abundances
We compared our NLTE atmospheric parameters and chemical abundances with previous studies that investigated the impact of NLTE analysis, focusing in particular on the behaviour of NLTE–LTE differences as a function of metallicity. Ruchti et al. (2013) introduced an optimised methodology, termed NLTE-Opt, specifically designed to mitigate the systematic biases typical of standard 1D LTE analyses for FGK-type stars. In this approach, is determined from Balmer lines profiles, while and [Fe/H] are derived from the NLTE ionization balance of Fe. Fig. 29 shows the systematic offsets in stellar parameters as a function of metallicity based on standard LTE analysis when compared with NLTE-Opt estimates. These trends are generally attributed to the inability of the LTE Fe I excitation balance to provide an accurate scale. As shown by the black circles representing the results of Ruchti et al. (2013), these differences decrease moving toward higher metallicities (see also Kovalev et al. 2019). Our Cepheids (in blue) follow the same general behaviour, displaying a good agreement with Ruchti et al. (2013), particularly in (panel a), (panel b) and (panel d). However, the panel c) of the same figure shows the difference in microturbulence differences, revealing an offset of 0.20–0.30 . Since Cepheids typically exhibit higher micro- and macroturbulent velocities, both the observed trends and the intrinsic scatter should be interpreted with caution. We also note that is the only parameter for which we find a statistically significant offset between the NLTE and LTE determinations (, see Fig. 2). Finally, our fitting method simultaneously optimizes all atmospheric parameters, which means that the contribution of significantly affects the determination of the others.
These results were obtained by varying all parameters simultaneously, performing the analysis once in NLTE and once in LTE. We further carried out an additional test to quantify the specific contribution of the NLTE versus LTE treatment to . In this test, we measured the NLTE and LTE metallicity of the dS23 sample while fixing , , and to their estimates. Fig. 30 shows the (NLTE–LTE) as a function of the LTE iron abundance, compared with the results of Ruchti et al. (2013). Our estimates closely follow the general behaviour reported by Ruchti et al. (2013) and remain well within 0.1 dex. These results further demonstrate that the offsets and trends observed in the comparison of metallicities with the literature (Figs. 3 and 4) are not primarily driven by the NLTE versus LTE treatment, but instead mainly reflect methodological differences, i.e. the adopted spectral analysis techniques and line lists.
@@
We independently verified our NLTE abundances by applying the line-by-line NLTE abundance corrections provided by the MPIA database161616https://nlte.mpia.de/ (Kovalev et al. 2018). Specifically, we focused on Mn and O. The atomic model for Mn is the same as that used in our study from Bergemann et al. (2019), while for O, we used the atomic model from Amarsi et al. (2018), while the MPIA database relies on Bergemann et al. (2021). We performed this comparison for five sample spectra whose atmospheric parameters are representative of the whole population. We downloaded the NLTE abundance correction from the MPIA database for Mn (5394.68 , 5420.33 and 6021.77 ) and O (7771.94 , 7774.17 and 7775.39 ). As shown in Fig. 31, the comparison indicates good agreement for all Cepheids, especially for Mn. For oxygen the agreement is also good (differences within 0.1 dex) for three Cepheids, while the remaining two show discrepancies of approximately 0.2 dex. This is well within our observational uncertainties. Finally, the decreasing trend observed in the plane vs has also been obtained in Vasilyev et al. (2019) for the 777.4 nm line with both 1D and 2D models, and previously observed in Cepheids in Luck et al. (2013).